### 15 Mar 2026 v2

# rank-33 generalized Clifford manifold and its twistor space

Abstract.
We introduce the notion of a rank-3 generalized Clifford manifold, defined by a triple of generalized complex structures satisfying Clifford-type relations. We show that every such structure canonically induces a generalized hypercomplex structure. We further describe a natural Spin(3)-action by Clifford rotations, which produces an $S^{2}\times S^{2}$-family of generalized complex structures. The corresponding twistor space is then constructed, and we prove that the induced almost generalized complex structure is integrable. In contrast to the standard pure-spinor approach, the integrability of the twistor-space structure is established entirely in terms of the generalized Nijenhuis tensor. We further prove that this Clifford-to-twistor construction is compatible with T-duality, in the sense that T-duality preserves the rank-3 Clifford triple, the induced structures, and the associated Spin(3)-rotated family.
The first author is supported by National Nature Science Foundation in China (Nos. 12101564) and Scientific Research Fund of Zhejiang Provincial Education Department (No. Y202456224). The second author is supported by Natural Science Foundation of Zhejiang Province (No. LMS26A010006). The third author is supported by Taishan Scholars Program for Young Experts of Shandong Province (No. tsqn202507265).
The authors would like to express their sincere gratitude to Professor Hanfei for his detailed and enlightening introduction to generalized complex geometry and T-duality. They are also deeply grateful to Professor Wei Wang for his illuminating exposition of classical twistor theory and Clifford structures. Their generous guidance and valuable insights have played an important role in the development of this work.
1. Introduction
The development of generalized geometric structures has profoundly influenced modern differential geometry, particularly through its deep connections with string theory and supersymmetric sigma models. The foundational concept of generalized complex geometry was introduced by Nigel Hitchin [14] and later systematically developed by Marco Gualtieri [11]. By replacing the tangent bundle $TM$ with the generalized tangent bundle $TM\oplus T^{*}M$, equipped with the Courant bracket and a natural neutral pairing, this theory provides a unified framework encompassing both complex and symplectic geometry. Within this framework, generalized Kähler geometry is defined by a pair of commuting generalized complex structures that are compatible with a common generalized metric (see [11, 12]). This structure extends classical Kähler geometry and naturally appears in two-dimensional supersymmetric sigma models with torsion.
A further step in this direction is the extension toward hyperkähler-type structures. In the classical setting, hyperkähler geometry is governed by three complex structures satisfying quaternionic relations. Motivated by this, Andreas Bredthauer [4] introduced generalized hyperkähler geometry by considering triples of generalized complex structures satisfying quaternionic-type algebraic relations. This framework allows one to describe supersymmetric backgrounds beyond the classical hyperkähler setting [7]. Later, Kimura, Sasaki, and Shiozawa [15] proposed a broader formulation of generalized hyperkähler structures based on the algebra of split-biquaternions. Their approach unifies aspects of hyperkähler and bi-hypercomplex geometries and naturally appears in $\mathcal{N}=(4,4)$ supersymmetric nonlinear sigma models. More recently, Fino and Grantcharov [9] studied generalized hypercomplex structures as $S^{2}$-families of anti-commuting generalized complex structures. Moreover, further discussions on generalized twistor spaces of hypercomplex structures can be found in [6] and [10].
From a structural viewpoint, hypercomplex geometry on Courant algebroids was systematically formulated by Stiénon [17]. In that work, hypercomplex structures are defined in terms of triples of orthogonal endomorphisms satisfying quaternionic relations, and it is proved that the vanishing of the associated Nijenhuis concomitants is equivalent to the existence of an (almost) torsion-free hypercomplex connection preserving the three complex structures. This result establishes a precise correspondence between integrability conditions and connection-theoretic characterizations. In [13], Wei Hong and Mathieu Stiénon proved that holomorphic symplectic structures and hypercomplex structures on Courant algebroids are equivalent notions.
More recently, hypercomplex geometry has been further broadened to include configurations in which the constituent complex structures satisfy Clifford-type rather than strictly quaternionic algebraic relations [16]. Such structures are closely related to bi-hypercomplex and bi-HKT (hyperkähler with torsion) geometries and naturally arise in models involving both ordinary and mirror multiplets [8]. The associated Clifford algebras often decompose into direct sums of quaternionic components, reflecting a bisector-type geometric structure and connecting to the doubled target-space formalism in string theory.
In this paper, we introduce the notion of a rank-$r$ almost generalized Clifford structure, which consists of $r$ almost generalized complex structures
$\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{r}$
satisfying the Clifford-type relations
$\mathcal{I}_{i}\mathcal{I}_{j}+\mathcal{I}_{j}\mathcal{I}_{i}=-2\delta_{ij}\mathrm{Id},\qquad i,j=1,\dots,r.$
When each $\mathcal{I}_{i}$ is integrable, we call such a structure a rank-$r$ generalized Clifford structure.
In what follows, we focus exclusively on the rank-3 case. Unless explicitly stated otherwise, the term generalized Clifford manifold will always mean a rank-3 generalized Clifford manifold. We will use the techniques in [8], [15] and [4] to study the generalized Clifford structure. By the equivalence of the Clifford algebra $Cl_{2,1}(\mathbb{R})$ and bi-quaternions, take
(1.1) $\mathcal{J}_{i}:=\frac{1}{2}\epsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k},\quad\mathcal{G}:=-\mathcal{I}_{1}\mathcal{J}_{1}=-\mathcal{I}_{2}\mathcal{J}_{2}=-\mathcal{I}_{3}\mathcal{J}_{3}=-\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}.$
We have
(1.2) $\begin{array}[]{ll}\mathcal{I}_{i}\mathcal{I}_{j}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{J}_{k},&\mathcal{J}_{i}\mathcal{J}_{j}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{J}_{k},\\ \mathcal{I}_{i}\mathcal{J}_{j}=-\delta_{ij}\mathcal{G}+\epsilon_{ijk}\mathcal{I}_{k},&\mathcal{J}_{i}\mathcal{I}_{j}=-\delta_{ij}\mathcal{G}+\epsilon_{ijk}\mathcal{I}_{k}\end{array}$
and
$\mathcal{G}^{2}=1.$
Our first main result shows that, in the rank-$3$ case, the individual integrability of the generalized Clifford generators already forces the simultaneous integrability of the whole induced generalized hypercomplex structure.
Theorem 1.1.
Let $\left(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ be an almost generalized Clifford structure on $\mathbb{T}M:=TM\oplus T^{*}M$, then
(1.3) $N\left(\mathcal{I}_{i},\mathcal{I}_{i}\right)=0\ (i=1,2,3)$
is equivalent to
(1.4) $N\left(\mathcal{I}_{i},\mathcal{I}_{j}\right)=N\left(\mathcal{J}_{i},\mathcal{J}_{j}\right)=N\left(\mathcal{I}_{i},\mathcal{J}_{j}\right)=0\ (i,j=1,2,3).$
where $\mathcal{J}_{i}$ is the induced almost generalized complex structure defined by (1.1).
We further show that rank-$3$ generalized Clifford structures admit natural $Spin(3)$-transformations induced by orthogonal rotations. These transformations act as rank-$3$ Clifford rotations and generate an $S^{2}\times S^{2}$-family of generalized complex structures, extending the classical twistor sphere construction. Via stereographic projection, each Riemann sphere is identified with $S^{2}\subset\mathbb{R}^{3}$. The associated rotation matrices $T\left(\zeta_{1}\right)$ and $S\left(\zeta_{2}\right)\in SO(3)$ encode the Clifford rotations determined by the parameters $\zeta_{1}$ and $\zeta_{2}$, which are the stereographic complex coordinates on the two factors of $S^{2}\times S^{2}$. The explicit formulas for these matrices are provided in Section 4. So we establish a twistor transformation for generalized Clifford manifolds, showing that Clifford rotations generate a natural $S^{2}\times S^{2}$-family of generalized complex structures over $TM\oplus T^{*}M$.
Theorem 1.2.
The transform
$\left(\begin{array}[]{c}\mathcal{I}_{1}\\ \mathcal{I}_{2}\\ \mathcal{I}_{3}\end{array}\right)\rightarrow\left(\begin{array}[]{c}\mathcal{K}_{1}(\zeta_{1},\zeta_{2})\\ \mathcal{K}_{2}(\zeta_{1},\zeta_{2})\\ \mathcal{K}_{3}(\zeta_{1},\zeta_{2})\end{array}\right)=\frac{1}{2}T(\zeta_{1})\left(\begin{array}[]{c}\mathcal{I}_{1}+\mathcal{I}_{2}\mathcal{I}_{3}\\ \mathcal{I}_{2}+\mathcal{I}_{3}\mathcal{I}_{1}\\ \mathcal{I}_{3}+\mathcal{I}_{1}\mathcal{I}_{2}\end{array}\right)+\frac{1}{2}S(\zeta_{2})\left(\begin{array}[]{c}-\mathcal{I}_{1}+\mathcal{I}_{2}\mathcal{I}_{3}\\ -\mathcal{I}_{2}+\mathcal{I}_{3}\mathcal{I}_{1}\\ -\mathcal{I}_{3}+\mathcal{I}_{1}\mathcal{I}_{2}\end{array}\right)$
defines a rank-3 Clifford rotation. Moreover, the triple
$\left(\mathcal{K}_{1}(\zeta_{1},\zeta_{2}),\mathcal{K}_{2}(\zeta_{1},\zeta_{2}),\mathcal{K}_{3}(\zeta_{1},\zeta_{2})\right)$
forms a new rank-$3$ generalized Clifford structure.
Furthermore, by Theorem 1.2, and in close analogy with the generalized hyperkähler case [4], we construct the twistor space of a rank-$3$ generalized Clifford manifold. We then prove, in terms of the generalized Nijenhuis tensor, that the induced generalized complex structure on the twistor space is integrable.
Theorem 1.3.
The twistor space of a rank-$3$ generalized Clifford manifold $M$ is the smooth manifold $Z=M\times S^{2}\times S^{2}$ endowed with the generalized complex structure $\mathbb{\hat{I}}$ defined at $T_{(p,\zeta_{1},\zeta_{2})}Z\oplus T_{(p,\zeta_{1},\zeta_{2})}^{*}Z$ as
$\left.\mathbb{\hat{I}}\right|_{(p,\zeta_{1},\zeta_{2})}=\left.\left.\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})\right|_{T_{p}M\oplus T_{p}^{*}M}\oplus\mathcal{J}\right|_{T\left(S^{2}\times S^{2}\right)\oplus T^{*}\left(S^{2}\times S^{2}\right)},$
where $\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})=\mathcal{K}_{1}(\zeta_{1},\zeta_{2})$ and $\mathcal{J}:=\left(\begin{array}[]{cccc}-J_{\zeta_{1}}&0&0&0\\ 0&-J_{\zeta_{2}}&0&0\\ 0&0&J^{T}_{\zeta_{1}}&0\\ 0&0&0&J^{T}_{\zeta_{2}}\end{array}\right)$ is the generalized complex structure on $S^{2}\times S^{2}$ induced by the canonical complex structure via $S^{2}\cong\mathbb{C}P^{1}$.
In Sect. 2, we investigate several structural properties of generalized complex structures in terms of the generalized Nijenhuis tensor. These results play a crucial role in proving the integrability of the generalized complex structure on the twistor space. In Sect. 3, we introduce generalized Clifford structures and present several basic examples. In Sect. 4, we construct the Spin(3)-transformations associated with rank-3 generalized Clifford structures and establish the generalized complex structure on the twistor space. In Sect. 5, we further study the behavior of rank-3 twisted generalized Clifford structures under T-duality, showing that T-duality preserves not only the original Clifford triple, but also the induced structures and the associated Spin(3)-rotated family.
2. Preliminaries
Let $M$ be a smooth manifold of even dimension $n$. Let $\mathbb{T}M=TM\oplus T^{*}M$, which is a direct sum of tangent bundle and cotangent bundle of $M$. The inner product $\langle\cdot,\cdot\rangle$ over $\mathbb{T}M$ is defined by
$\langle X+\xi,Y+\eta\rangle:=\frac{1}{2}(\xi(Y)+\eta(X)),$
where $X,Y\in TM$ and $\xi,\eta\in TM^{*}$. Obviously, the inner product $\langle\cdot,\cdot\rangle$ is symmetric and has signature $(n,n)$. The Dorfman bracket $\cdot,\cdot]_{D}$ over $\mathbb{T}M$ is defined by
(2.1) $X+\xi,Y+\eta]_{D}=[X,Y]+L_{X}\eta-\iota_{Y}{\rm d}\xi.$
The Dorfman bracket satisfies the Jacobi identity but is not skew-symmetric. The twisted version of a bracket $X+\xi,Y+\eta]_{D}$ over $\mathbb{T}M$ with a $3$-form $H$ is defined by
(2.2) $X+\xi,Y+\eta]_{H}=[X,Y]+L_{X}\eta-\iota_{Y}{\rm d}\xi-\iota_{Y}\iota_{X}H.$
The Jacobi identity for the bracket $X+\xi,Y+\eta]_{H}$ requires that H must be closed.
2.1. Generalized complex structure
In this section, we formulate several properties of generalized complex structures purely in terms of the generalized Nijenhuis tensor, which will be used in the proofs of the subsequent theorems.
An almost generalized complex structure on $M$ is an orthogonal endomorphism $\mathcal{J}$ of $\mathbb{T}M$ satisfying $\mathcal{J}^{2}=$- Id. If $\mathcal{J}$ is integrable with respect to the Dorfman bracket, then it is called a generalized complex structure. For two bundle endomorphisms
$\mathcal{I},\mathcal{J}:\mathbb{T}M\rightarrow\mathbb{T}M$
their mixed Nijenhuis tensor (also called the Nijenhuis concomitant) is the bilinear map
$\mathcal{N}(\mathcal{I},\mathcal{J}):\Gamma(\mathbb{T}M)\times\Gamma(\mathbb{T}M)\rightarrow\Gamma(\mathbb{T}M)$
defined by
(2.3) $\mathcal{N}(\mathcal{I},\mathcal{J})\left(A,B\right)=$ $\frac{1}{2}\left({\left[\mathcal{I}A,\mathcal{J}B\right]_{D}+\left[\mathcal{J}A,\mathcal{I}B\right]_{D}}-\mathcal{I}\left[A,\mathcal{J}B\right]_{D}-\mathcal{I}\left[\mathcal{J}A,B\right]_{D}\right$.
$\left.-\mathcal{J}\left[A,\mathcal{I}B\right]_{D}-\mathcal{J}\left[\mathcal{I}A,B\right]_{D}+\mathcal{I}\mathcal{J}\left[A,B\right]_{D}+\mathcal{J}\mathcal{I}\left[A,B\right]_{D}\right),$
where $\cdot,\cdot]_{D}$ is the Dorfman bracket and $A,B\in\Gamma(\mathbb{T}M)$. And the Nijenhuis tensor of $\mathcal{J}$ is
(2.4) $N_{\mathcal{J}}(A,B)=\mathcal{N}(\mathcal{J},\mathcal{J})\left(A,B\right)=[\mathcal{J}A,\mathcal{J}B]_{D}-\mathcal{J}[\mathcal{J}A,B]_{D}-\mathcal{J}[A,\mathcal{J}B]_{D}-[A,B]_{D}.$
Lemma 2.1.
For generalized almost-complex structure $\mathcal{I}$ and $\mathcal{J}$, if $\mathcal{I}\mathcal{J}=-\mathcal{J}\mathcal{I}$, we have
  • (1)
    The Nijenhuis expression for $\mathcal{I}\mathcal{J}$:
    $\begin{split}N_{\mathcal{I}\mathcal{J}}(A,B)=&\frac{1}{2}\left[N_{\mathcal{I}}(\mathcal{J}A,\mathcal{J}B)+N_{\mathcal{J}}(\mathcal{I}A,\mathcal{I}B)+N_{\mathcal{I}}(A,B)+N_{\mathcal{J}}(A,B)\right.\\ -&\left.\mathcal{I}N_{J}(\mathcal{I}A,B)-\mathcal{I}N_{\mathcal{J}}(A,\mathcal{I}B)-\mathcal{J}N_{\mathcal{I}}(\mathcal{J}A,B)-\mathcal{J}N_{\mathcal{I}}(A,\mathcal{J}B)\right].\end{split}$
  • (2)
    The mixed Nijenhuis expression for $\mathcal{I}$ and $\mathcal{J}$:
    $\begin{split}N\left(\mathcal{I},\mathcal{J}\right)(A,B)=&\frac{1}{2}\left[-N_{\mathcal{I}\mathcal{J}}(\mathcal{J}A,\mathcal{I}B)+\mathcal{I}\mathcal{J}\left(N_{\mathcal{I}}(A,B)-N_{\mathcal{J}}(A,B)\right)\right].\end{split}$
Proof.
(1) Obviously, with respect to the Dorfman bracket (2.1), we have
$\mathcal{I}\mathcal{J}A,B]_{D}=-[\mathcal{J}\mathcal{I}A,B]_{D}$
since $\mathcal{I}\mathcal{J}=-\mathcal{JI}$. And the Dorfman bracket is linear in its first argument, i.e.,
$A+B,C]_{D}=[A,C]_{D}+[B,C]_{D}.$
Then we have
(2.5) $\begin{split}&2N_{\mathcal{I}\mathcal{J}}(X,Y)\\ =&[\mathcal{I}\mathcal{J}A,\mathcal{I}\mathcal{J}B]_{D}-\mathcal{I}\mathcal{J}[\mathcal{I}\mathcal{J}A,B]_{D}-\mathcal{I}\mathcal{J}[A,\mathcal{I}\mathcal{J}B]_{D}-[A,B]_{D}\\ +&[\mathcal{J}\mathcal{I}A,\mathcal{J}\mathcal{I}B]_{D}-\mathcal{J}\mathcal{I}[\mathcal{J}\mathcal{I}A,B]_{D}-\mathcal{J}\mathcal{I}[A,\mathcal{J}\mathcal{I}B]_{D}-[A,B]_{D}\\ =&\left(N_{\mathcal{I}}(\mathcal{J}A,\mathcal{J}B)+{\mathcal{I}[\mathcal{I}\mathcal{J}A,\mathcal{J}B]_{D}}+\mathcal{I}[\mathcal{J}A,\mathcal{I}\mathcal{J}B]_{D}+[\mathcal{J}A,\mathcal{J}B]_{D}\right)-{\mathcal{I}\mathcal{J}[\mathcal{I}\mathcal{J}A,B]_{D}}-\mathcal{I}\mathcal{J}[A,\mathcal{I}\mathcal{J}B]_{D}-[A,B]_{D}\\ +&\left(N_{\mathcal{J}}(\mathcal{I}A,\mathcal{I}B)+{\mathcal{J}[\mathcal{J}\mathcal{I}A,\mathcal{I}B]_{D}}+\mathcal{J}[\mathcal{I}A,\mathcal{J}\mathcal{I}B]_{D}+[\mathcal{I}A,\mathcal{I}B]_{D}\right)-{\mathcal{J}\mathcal{I}[\mathcal{J}\mathcal{I}A,B]_{D}}-\mathcal{J}\mathcal{I}[A,\mathcal{J}\mathcal{I}B]_{D}-[A,B]_{D},\end{split}$
by the definition of $N_{\mathcal{I}}(\mathcal{J}A,\mathcal{J}B)$ and $N_{\mathcal{J}}(\mathcal{I}A,\mathcal{I}B)$. And we have,
[…]
by $\mathcal{I}\mathcal{J}=-\mathcal{JI}$ and the definition of $N_{\mathcal{J}}(\mathcal{I}A,B)$. Similarly,
[…]
Substituting (2.6)-(2.7) into (2.5) yields
(2.8) $\begin{split}&2N_{\mathcal{I}\mathcal{J}}(A,B)\\ =&N_{\mathcal{I}}(\mathcal{J}A,\mathcal{J}B)+N_{\mathcal{J}}(\mathcal{I}A,\mathcal{I}B)-\mathcal{I}N_{\mathcal{J}}(\mathcal{I}A,B)-\mathcal{I}N_{\mathcal{J}}(A,\mathcal{I}B)-\mathcal{J}N_{\mathcal{I}}(\mathcal{J}A,B)-\mathcal{J}N_{\mathcal{I}}(A,\mathcal{J}B)\\ +&N_{\mathcal{I}}(A,B)+N_{\mathcal{J}}(A,B),\end{split}$
by $\mathcal{I}\mathcal{J}=-\mathcal{JI}$ and the definition of Nijenhuis tensor.
(2) By the definition (2.3) and $\mathcal{I}\mathcal{J}=-\mathcal{J}\mathcal{I}$,
(2.9) $2\mathcal{N}(\mathcal{I},\mathcal{J})\left(A,B\right)=$ ${\left[\mathcal{I}A,\mathcal{J}B\right]_{D}+\left[\mathcal{J}A,\mathcal{I}B\right]_{D}}-\mathcal{I}\left[A,\mathcal{J}B\right]_{D}-\mathcal{I}\left[\mathcal{J}A,B\right]_{D}-\mathcal{J}\left[A,\mathcal{I}B\right]_{D}-\mathcal{J}\left[\mathcal{I}A,B\right]_{D}.$
By the definition of $\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)$, $\mathcal{N}_{\mathcal{I}}\left(A,B\right)$ and $\mathcal{N}_{\mathcal{J}}\left(A,B\right)$, we have
(2.10) $\begin{split}&{\left[\mathcal{I}A,\mathcal{J}B\right]_{D}+\left[\mathcal{J}A,\mathcal{I}B\right]_{D}}\\ =&-\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\left[\mathcal{I}A,\mathcal{I}B\right]_{D}-\left[\mathcal{J}A,\mathcal{J}B\right]_{D}\right)\\ =&-\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{I}}\left(A,B\right)+\mathcal{I}\left[\mathcal{I}A,B\right]_{D}+\mathcal{I}\left[A,\mathcal{I}B\right]_{D}+\left[A,B\right]_{D}\right)\\ &-\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{J}}\left(A,B\right)+\mathcal{J}\left[\mathcal{J}A,B\right]_{D}+\mathcal{J}\left[A,\mathcal{J}B\right]_{D}+\left[A,B\right]_{D}\right)\\ =&-\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{I}}\left(A,B\right)-\mathcal{N}_{\mathcal{J}}\left(A,B\right)\right)+\mathcal{J}\left[\mathcal{I}A,B\right]_{D}+\mathcal{J}\left[A,\mathcal{I}B\right]_{D}+\mathcal{I}\left[\mathcal{J}A,B\right]_{D}+\mathcal{I}\left[A,\mathcal{J}B\right]_{D}.\end{split}$
Substituting (2.10) into (2.9) yields
$2\mathcal{N}(\mathcal{I},\mathcal{J})\left(A,B\right)=$ $-\mathcal{N}_{\mathcal{I}\mathcal{J}}\left(\mathcal{J}A,\mathcal{I}B\right)+\mathcal{I}\mathcal{J}\left(\mathcal{N}_{\mathcal{I}}\left(A,B\right)-\mathcal{N}_{\mathcal{J}}\left(A,B\right)\right),$
i.e., the second term of Lemma 2.1 holds. ∎
Then the following result in [17] and [13] is a direct corollary of the Lemma 2.1.
Lemma $2.$
([17, 13]). For generalized complex structures $\mathcal{I}$ and $\mathcal{J}$, if $\mathcal{I}\mathcal{J}=-\mathcal{J}\mathcal{I}$, then
  • (1)
    $N_{\mathcal{I}\mathcal{J}}=0$;
  • (2)
    $N\left(\mathcal{I},\mathcal{J}\right)=0$.
Proof.
(1) Because $N_{\mathcal{I}}=N_{\mathcal{J}}=0$, the conclusion (1) follows immediately from the first term of Lemma 2.1.
(2)By $N_{\mathcal{I}}=N_{\mathcal{J}}=0$, Lemma 2.1–(2) and Lemma 2.2–(1), (2) holds clearly.
Lemma 2.3.
For generalized complex structure $\mathcal{I}$ and $\mathcal{J}$, if $\mathcal{I}\mathcal{J}=\mathcal{J}\mathcal{I}$, we have $N(\mathcal{I},\mathcal{J})=0$.
Proof.
Since $\mathcal{I}$ and $\mathcal{J}$ commute and satisfy $\mathcal{I}^{2}=\mathcal{J}^{2}=-1$, they admit a simultaneous eigenspace decomposition of the complexified bundle
$\mathbb{T}M_{\mathbb{C}}=\bigoplus_{a,b=\pm 1}L^{a,b},\quad L^{a,b}=\left\{A\in\mathbb{T}M_{\mathbb{C}}\mid\mathcal{I}A=a\mathbf{i}A,\mathcal{J}A=b\mathbf{i}A\right\}.$
Integrability of $\mathcal{I}$ and $\mathcal{J}$ is equivalent to involutivity of their eigenbundles. Hence each joint eigenspace $L^{a,b}$ is closed under the Dorfman bracket. Let $A_{1},A_{2}\in\Gamma\left(L^{a,b}\right)$. Using $\mathcal{I}A_{k}=a\mathbf{i}A_{k},\mathcal{J}A_{k}=b\mathbf{i}A_{k}$ and involutivity of $L^{a,b}$, a direct substitution into the definition of the Nijenhuis concomitant (2.3) shows
$\mathcal{N}(\mathcal{I},\mathcal{J})\left(A_{1},A_{2}\right)=0.$
In fact,
(2.11) $\mathcal{N}(\mathcal{I},\mathcal{J})\left(A_{1},A_{2}\right)=$ $\frac{1}{2}\left({\left[\mathcal{I}A_{1},\mathcal{J}A_{2}\right]_{D}+\left[\mathcal{J}A_{1},\mathcal{I}A_{2}\right]_{D}}-\mathcal{I}\left[A_{1},\mathcal{J}A_{2}\right]_{D}-\mathcal{I}\left[\mathcal{J}A_{1},A_{2}\right]_{D}\right$.
$\left.-\mathcal{J}\left[A_{1},\mathcal{I}A_{2}\right]_{D}-\mathcal{J}\left[\mathcal{I}A_{1},A_{2}\right]_{D}+\mathcal{I}\mathcal{J}\left[A_{1},A_{2}\right]_{D}+\mathcal{J}\mathcal{I}\left[A_{1},A_{2}\right]_{D}\right)$
$=$ $\frac{1}{2}\left(-ab-ba+ab+ab+ba+ba-ab-ba\right)\left[A_{1},A_{2}\right]_{D}$
$=$ $0,$
by $\mathcal{I}\mathcal{J}=\mathcal{J}\mathcal{I}$. By the $\mathbb{R}$-bilinearity of $\mathcal{N}(\mathcal{I},\mathcal{J})$ and the decomposition $\mathbb{T}M_{\mathbb{C}}=\oplus L^{a,b},$ the tensor vanishes identically:
$\mathcal{N}(\mathcal{I},\mathcal{J})=0.$
Lemma 2.4.
Let $\mathcal{I},\mathcal{J}\in\Gamma(\mathrm{End}\,\mathbb{T}M)$ satisfy
(2.12) $\mathcal{I}^{2}=-\mathrm{Id},\qquad\mathcal{I}\mathcal{J}+\mathcal{J}\mathcal{I}=0.$
Then
(2.13) $2\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})(A,B)=\mathcal{N}_{\mathcal{I}}(\mathcal{J}A,B)+\mathcal{N}_{\mathcal{I}}(A,\mathcal{J}B)+2\mathcal{I}\mathcal{N}(\mathcal{I},\mathcal{J})(A,B),$
where $A,B\in\Gamma(\mathbb{T}M)$.
Proof.
It follows from $\mathcal{I}^{2}=-\mathrm{Id}$ and (2.4) that
(2.14) $\mathcal{I}A,\mathcal{I}B]_{D}=\mathcal{N}_{\mathcal{I}}(A,B)+\mathcal{I}[\mathcal{I}A,B]_{D}+\mathcal{I}[A,\mathcal{I}B]_{D}+[A,B]_{D},$
for all $A,B\in\Gamma(\mathbb{T}M)$. Expanding $\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})$ via the definition (2.3), we obtain
(2.15) $2\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})(A,B)$ $=[\mathcal{I}A,\mathcal{I}\mathcal{J}B]_{D}+[\mathcal{I}\mathcal{J}A,\mathcal{I}B]_{D}-\mathcal{I}[A,\mathcal{I}\mathcal{J}B]_{D}-\mathcal{I}[\mathcal{I}\mathcal{J}A,B]_{D}$
$-(\mathcal{I}\mathcal{J})[A,\mathcal{I}B]_{D}-(\mathcal{I}\mathcal{J})[\mathcal{I}A,B]_{D}+\mathcal{I}(\mathcal{I}\mathcal{J})[A,B]_{D}+(\mathcal{I}\mathcal{J})\mathcal{I}[A,B]_{D}.$
Using $\mathcal{I}^{2}=-\mathrm{Id}$ and $\mathcal{I}\mathcal{J}=-\mathcal{J}\mathcal{I}$, we obtain
$\mathcal{I}(\mathcal{I}\mathcal{J})=-\mathcal{J},\qquad(\mathcal{I}\mathcal{J})\mathcal{I}=\mathcal{J},$
so the last two terms in (2.15) cancel. Hence
(2.16) $2\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})(A,B)$ $=[\mathcal{I}A,\mathcal{I}\mathcal{J}B]_{D}+[\mathcal{I}\mathcal{J}A,\mathcal{I}B]_{D}-\mathcal{I}\left([A,\mathcal{I}\mathcal{J}B]_{D}+[\mathcal{I}\mathcal{J}A,B]_{D}\right)-\mathcal{I}\mathcal{J}\left([A,\mathcal{I}B]_{D}+[\mathcal{I}A,B]_{D}\right).$
Since $\mathcal{I}\mathcal{J}B=\mathcal{I}(\mathcal{J}B)$ and $\mathcal{I}\mathcal{J}A=\mathcal{I}(\mathcal{J}A)$, apply (2.14) twice:
(2.17) $\begin{split}[\mathcal{I}A,\mathcal{I}(\mathcal{J}B)]_{D}&=\mathcal{N}_{\mathcal{I}}(A,\mathcal{J}B)+\mathcal{I}[\mathcal{I}A,\mathcal{J}B]_{D}+\mathcal{I}[A,\mathcal{I}\mathcal{J}B]_{D}+[A,\mathcal{J}B]_{D},\\ [\mathcal{I}(\mathcal{J}A),\mathcal{I}B]_{D}&=\mathcal{N}_{\mathcal{I}}(\mathcal{J}A,B)+\mathcal{I}[\mathcal{I}\mathcal{J}A,B]_{D}+\mathcal{I}[\mathcal{J}A,\mathcal{I}B]_{D}+[\mathcal{J}A,B]_{D}.\end{split}$
Substituting (2.17) into (2.16) cancels the middle $\mathcal{I}$-terms and gives
(2.18) $2\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})(A,B)$ $=\mathcal{N}_{\mathcal{I}}(\mathcal{J}A,B)+\mathcal{N}_{\mathcal{I}}(A,\mathcal{J}B)+\mathcal{I}[\mathcal{I}A,\mathcal{J}B]_{D}+\mathcal{I}[\mathcal{J}A,\mathcal{I}B]_{D}+[A,\mathcal{J}B]_{D}+[\mathcal{J}A,B]_{D}$
$\qquad-(\mathcal{I}\mathcal{J})[A,\mathcal{I}B]_{D}-(\mathcal{I}\mathcal{J})[\mathcal{I}A,B]_{D}.$
By the definition of $\mathcal{N}(\mathcal{I},\mathcal{J})$, we have
(2.19) $\mathcal{I}A,\mathcal{J}B]_{D}+[\mathcal{J}A,\mathcal{I}B]_{D}=2\mathcal{N}(\mathcal{I},\mathcal{J})(A,B)+\mathcal{I}[A,\mathcal{J}B]_{D}+\mathcal{I}[\mathcal{J}A,B]_{D}+\mathcal{J}[A,\mathcal{I}B]_{D}+\mathcal{J}[\mathcal{I}A,B]_{D}.$
Apply $\mathcal{I}$ to (2.19) and use $\mathcal{I}^{2}=-\mathrm{Id}$:
(2.20) $\mathcal{I}[\mathcal{I}A,\mathcal{J}B]_{D}+\mathcal{I}[\mathcal{J}A,\mathcal{I}B]_{D}$ $=2\mathcal{I}\mathcal{N}(\mathcal{I},\mathcal{J})(A,B)-[A,\mathcal{J}B]_{D}-[\mathcal{J}A,B]_{D}+(\mathcal{I}\mathcal{J})[A,\mathcal{I}B]_{D}+(\mathcal{I}\mathcal{J})[\mathcal{I}A,B]_{D}.$
Substituting (2.20) into (2.18) shows
$2\mathcal{N}(\mathcal{I},\mathcal{I}\mathcal{J})(A,B)=\mathcal{N}_{\mathcal{I}}(\mathcal{J}A,B)+\mathcal{N}_{\mathcal{I}}(A,\mathcal{J}B)+2\mathcal{I}\mathcal{N}(\mathcal{I},\mathcal{J})(A,B),$
proving the claim. ∎
2.2. Generalized real structure
Motivated by the geometric meaning of integrability for generalized complex structures, we introduce generalized real structures on a Dorfman algebroid and investigate their integrability. While the condition $\mathcal{G}^{2}=1$ is purely algebraic and only yields a splitting $\mathbb{T}M=\mathbb{T}M_{+}\oplus\mathbb{T}M_{-}$ into the $\pm 1$ eigenbundles, integrability upgrades this splitting to a genuine geometric decomposition: the eigenbundles become involutive, hence define Lie algebroid structures (and, under suitable isotropy assumptions, a pair of complementary Dirac structures).
Definition 2.1.
We call $\mathcal{G}$ a generalized almost-real structure if $\mathcal{G}^{2}=1$ and $\mathcal{G}\in O(n,n)$.
So the eigenspace of $\mathcal{G}$ w.r.t eigenvalue $1$ has the form
$Ker(\mathcal{G}-{\rm Id})=\{A+\mathcal{G}A|\forall A\in\mathbb{T}M\}$
and eigenspace of $\mathcal{G}$ w.r.t eigenvalue 1-1 has the form
$Ker(\mathcal{G}+{\rm Id})=\{A-\mathcal{G}A|\forall A\in\mathbb{T}M\}.$
We will show that the integrability is characterized by the vanishing of an associated Nijenhuis tensor. This provides a real analogue of generalized complex geometry and yields a natural framework for decomposing Dorfman algebroids into pairs of complementary Dirac structures.
Definition 2.2.
We call the generalized almost-real structure $\mathcal{G}$ a generalized real structure if $Ker(\mathcal{G}\pm{\rm Id})$ is closed with respect to the Dorfman bracket (2.1), i.e.
(2.21) $Ker(\mathcal{G}\pm{\rm Id}),Ker(\mathcal{G}\pm{\rm Id})]\subseteq Ker(\mathcal{G}\pm{\rm Id}).$
We define the Nijenhuis tensor of real structure $\mathcal{G}$ as
(2.22) $N_{\mathcal{G}}(A,B)=[\mathcal{G}A,\mathcal{G}B]_{D}-\mathcal{G}[\mathcal{G}A,B]_{D}-\mathcal{G}[A,\mathcal{G}B]_{D}+[A,B]_{D}\quad(A,B\in\mathbb{T}M),$
which differs slightly from the definition of the Nijenhuis tensor in the case of a generalized complex structure.
Lemma 2.5.
The vanishing of $N_{\mathcal{G}}$ is equivalent to the integrability of the real structure $\mathcal{G}$.
Proof.
Obviously, for $A,B\in\mathbb{T}M$, we have
(2.23) $A\pm\mathcal{G}A,B\pm\mathcal{G}B]_{D}=[A,B]_{D}\pm[A,\mathcal{G}B]_{D}\pm[\mathcal{G}A,B]_{D}+[\mathcal{G}A,\mathcal{G}B]_{D}.$
If $\mathcal{G}\left([A,\mathcal{G}B]_{D}+[\mathcal{G}A,B]_{D}\right)=[A,B]_{D}+[\mathcal{G}A,\mathcal{G}B]_{D},$ i.e. $N_{\mathcal{G}}(A,B)=0,$ we have
$\mathcal{G}[A\pm\mathcal{G}A,B\pm\mathcal{G}B]_{D}=\pm[A\pm\mathcal{G}A,B\pm\mathcal{G}B]_{D},$
i.e. $\mathcal{G}$ is integrable. ∎
The generalized metric, introduced by Hitchin and developed by Gualtieri, is the analogue of a Riemannian metric in generalized geometry. A generalized metric is an endomorphism $\mathcal{G}\in\mathrm{End}(\mathbb{T}M)\cap O(n,n)$ such that $\mathcal{G}^{2}=\mathrm{Id}$, and $\langle\mathcal{G}A,A\rangle>0$ for all nonzero $A\in\mathbb{T}M$. And generalized metric is equivalent to pairs $(g,b)$ consisting of a Riemannian metric $g$ and a $2$-form $b$. And it has the specific expression
$\mathcal{G}=\left(\begin{array}[]{ll}1&0\\ b&1\end{array}\right)\left(\begin{array}[]{cc}0&g^{-1}\\ g&0\end{array}\right)\left(\begin{array}[]{cc}1&0\\ -b&1\end{array}\right)=\left(\begin{array}[]{cc}-g^{-1}b&g^{-1}\\ g-bg^{-1}b&bg^{-1}\end{array}\right).$
However, it is not integrable.
Proposition 2.1.
The generalized metric $\mathcal{G}$, viewed as a generalized almost real structure, is in general not integrable.
Proof.
Let $\mathcal{G}$ be the generalized metric determined by a Riemannian metric $g$ and a 2 -form $b$. Its +1 eigenbundle is $C_{+}=\{X+(b+g)(X)\mid X\in TM\}.$ The Dorfman bracket (2.1) satisfies the compatibility condition
$A,A]_{D}=\mathcal{D}\langle A,A\rangle,$
where $\mathcal{D}$ is the Courant algebroid differential associated with the anchor map $\rho$, defined by $\langle\mathcal{D}f,A\rangle=\frac{1}{2}\rho(A)f$. Taking $A=X+(b+g)(X)\in C_{+}$, we have
$A,A]_{D}=\mathcal{D}\langle X+(b+g)(X),X+(b+g)(X)\rangle=\mathcal{D}(g(X,X)).$
If $\mathcal{G}$ were an integrable generalized real structure, then $C_{+}$ would be closed under the Dorfman bracket, hence
$A,A]_{D}\in C_{+}\quad\text{ for all }A\in C_{+}.$
But $\mathcal{D}(g(X,X))$ is a pure 1-form, while $C_{+}\cap T^{*}M=\{0\}$. Therefore one must have
$\mathcal{D}(g(X,X))=0,$
or equivalently,
${\rm d}(g(X,X))=0\quad\text{ for all }X\in TM.$
This condition is incompatible with a nondegenerate metric $g$ in general. Hence the eigenbundles $C_{\pm}$ of a generalized metric are not closed under the Dorfman bracket, and the associated almost real structure $\mathcal{G}$ is typically not integrable. ∎
3. Generalized Clifford manifold
A rank-3 Clifford-Hermitian manifold is a Riemannian manifold whose tangent bundle carries a pointwise representation of the rank-3 Clifford algebra [16]. Concretely, it is equipped with three anticommuting orthogonal almost complex structures $I_{1},I_{2},I_{3}$, each compatible with the Riemannian metric. As a natural generalization, we introduce the generalized Clifford manifold by generalized complex structures on the Dorfman algebroid $TM\oplus T^{*}M$. This extension preserves many structural features of the classical theory while greatly enlarging the class of admissible geometries.
Definition 3.1.
A (twisted) rank-$r$ generalized Clifford structure on $\mathbb{T}M$ consists of $r$ almost generalized complex structures
$I_{1},\dots,I_{r}\in\mathrm{End}(\mathbb{T}M)$
satisfying the Clifford-type relations
(3.1) $\mathcal{I}_{i}\mathcal{I}_{j}+\mathcal{I}_{j}\mathcal{I}_{i}=-2\delta_{ij}\mathrm{Id},$
together with the integrability conditions
(3.2) $\mathcal{N}(\mathcal{I}_{i},\mathcal{I}_{i})=0,\quad i=1,2,\cdots,r,$
where $\mathcal{N}(\cdot,\cdot)$ denotes the generalized Nijenhuis concomitant associated with the (twisted) Dorfman bracket. A manifold endowed with such a structure is called a (twisted) rank-$r$ generalized Clifford manifold.
In the rank-$3$ case, Theorem 1.1 shows that the mixed generalized Nijenhuis concomitants vanish automatically, so that every rank-$3$ generalized Clifford structure canonically determines a generalized hypercomplex structure.
Proof of Theorem 1.1 ($\Longrightarrow$) When $i\neq j$, substituting $\mathcal{I}=\mathcal{I}_{i}$ and $\mathcal{J}=\mathcal{I}_{j}$ (with $i\neq j$ ) into the second part of Lemma 2.2, we obtain
(3.3) $\mathcal{N}\left(\mathcal{I}_{i},\mathcal{I}_{j}\right)=0.$
Therefore, (3.2) and (3.3) imply
(3.4) $N\left(\mathcal{I}_{i},\mathcal{I}_{j}\right)=0\ (i,j=1,2,3).$
By substituting $\mathcal{I}=\mathcal{I}_{i}$ and $\mathcal{J}=\mathcal{I}_{j}\ (i\neq j)$ into the first term of Lemma 2.2, we obtain
(3.5) $\mathcal{N}\left(\mathcal{J}_{l},\mathcal{J}_{l}\right)=0,\ l=1,2,3,$
since $\mathcal{J}_{i}=\frac{1}{2}\epsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}$. Applying Lemma 2.2–(2) with $\mathcal{I}=\mathcal{J}_{i}$ and $\mathcal{J}=\mathcal{J}_{j}\ (i\neq j)$ yields
(3.6) $\mathcal{N}\left(\mathcal{J}_{l},\mathcal{J}_{m}\right)=0,\ l\neq m,$
by (3.5). Then (3.5) and (3.6) together imply
(3.7) $N\left(\mathcal{J}_{i},\mathcal{J}_{j}\right)=0\ (i,j=1,2,3).$
By identity (1.2), we have $\mathcal{I}_{i}\mathcal{J}_{i}=\mathcal{J}_{i}\mathcal{I}_{i}\ (i=1,2,3).$ Then we have
(3.8) $\mathcal{N}\left(\mathcal{I}_{i},\mathcal{J}_{i}\right)=0,\ i=1,2,3,$
by (3.2), (3.5) and Lemma 2.3. Applying Lemma 2.4 with $\mathcal{I}=\mathcal{I}_{i}$ and $\mathcal{J}=\mathcal{I}_{j}\ (i\neq j)$ yields
(3.9) $\mathcal{N}\left(\mathcal{I}_{l},\mathcal{J}_{m}\right)=0,\ l\neq m,$
by (3.2) and (3.3). It follows from (3.8) and (3.9)
(3.10) $N\left(\mathcal{I}_{i},\mathcal{J}_{j}\right)=0\ (i,j=1,2,3).$
By (3.4), (3.7) and (3.10), (1.4) holds.
[…]
A generalized hypercomplex structure [9] on $M$ is a triple $(\mathcal{I},\mathcal{J},\mathcal{K})$ of generalized complex structures on $\mathbb{T}M$ satisfying the quaternionic algebra: $\mathcal{I}^{2}=\mathcal{J}^{2}=\mathcal{K}^{2}=\mathcal{I}\mathcal{J}\mathcal{K}=-\mathrm{Id}.$ If each $\mathcal{I},\mathcal{J},\mathcal{K}$ is integrable in the generalized sense (i.e., its eigenbundle is Dorfman involutive), then $(M,\mathcal{I},\mathcal{J},\mathcal{K})$ is called a generalized hypercomplex manifold. Then by Lemma 2.2–(1), we have
Proposition 3.1.
The induced almost generalized complex structure $\mathcal{J}_{i}=\frac{1}{2}\epsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}$ in (1.1) is integrable. Then the generalized Clifford manifold canonically determines a generalized hypercomplex structure $\left(\mathcal{J}_{1},\mathcal{J}_{2},\mathcal{J}_{3}\right)$.
Example 3.1.
(Clifford–Hermitian manifolds). The rank-$3$ Clifford Hermitian manifold [16] is a Hermite manifold with three complex structure ${\rm I}_{a}$ $(a=1,2,3)$ such that
  • (1)
    $g({\rm I}_{a}X,{\rm I}_{a}Y)=g(X,Y)$, $X,Y\in TM$;
  • (2)
    $I_{a}I_{b}+I_{b}I_{a}=-2\delta_{ab}{\rm Id}.$
If the Clifford structures $I_{i}(i=1,2,3)$ are simultaneously integrable, that is,
$\mathcal{N}\left(I_{i},I_{j}\right)=0,\quad i,j=1,2,3,$
then the associated generalized complex structures
$\mathcal{I}_{i}=\left(\begin{array}[]{cc}I_{i}&0\\ 0&-I_{i}^{*}\end{array}\right),\quad i=1,2,3,$
define a generalized Clifford structure on the generalized tangent bundle $\mathbb{T}M=TM\oplus T^{*}M$.
Example 3.2.
(Generalized hyperkähler manifolds). Every generalized hyperkähler manifold gives rise to a generalized Clifford structure. This is because that the generalized hyperkähler manifold admits three generalized Kähler structures $\left(\mathcal{I}_{i},G\right)$ $(i=1,2,3)$ that satisfy the algebra of bi-quaternions $Cl_{2,1}(\mathbb{R})$
(3.11) $\begin{array}[]{ll}\mathcal{I}_{i}\mathcal{I}_{j}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{I}_{k},&\tilde{\mathcal{I}}_{i}{\tilde{\mathcal{I}}_{j}}=-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{I}_{k},\\ \mathcal{I}_{i}\tilde{\mathcal{I}}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},&\tilde{\mathcal{I}}_{i}\mathcal{I}_{j}=-\delta_{ij}{G}+\epsilon_{ijk}\tilde{\mathcal{I}}_{k},\end{array}$
where $\tilde{\mathcal{I}}_{i}=\mathcal{I}_{i}G$. It is direct to check that the three generalized complex structures $\tilde{\mathcal{I}}_{i}\ (i=1,2,3)$ form a generalized Clifford structure on $\mathbb{T}M$ .
Obviously, by (1.2) and (3.11), we have
Proposition 3.2.
The generalized Clifford manifold $\left(M,\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ induces a generalized hyperkähler structure when $\mathcal{G}=-\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}$ is a generalized metric.
Example 3.3.
(Hyperkähler manifolds). Every hyperkähler manifold yields a generalized Clifford structure in the generalized sense. In fact, take
$\mathcal{I}_{i}:=\left(\begin{array}[]{cc}0&-\omega_{I_{i}}^{-1}\\ \omega_{I_{i}}&0\end{array}\right),\quad i=1,2,3.$
It is direct to check that $\mathcal{I}_{i}\ (i=1,2,3)$ satisfies identity (3.1)-(3.2), i.e., $\left(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}\right)$ is a generalized Clifford structure.
Remark 3.1.
In the definition of a hyperkähler manifold with hypercomplex structure $I_{1},I_{2},I_{3}$, one does not need to impose the mixed Nijenhuis condition $N(I_{i},I_{j})=0\ (i\neq j)$ as an additional axiom. Indeed, for a hyperkähler structure the complex structures $I_{1},I_{2},I_{3}$ are parallel with respect to the Levi-Civita connection, hence each of them is integrable. As a consequence, any mixed Nijenhuis concomitant vanishes automatically. Therefore the condition $N(I_{i},I_{j})=0\ (i\neq j)$ is redundant in the hyperkähler setting.
Example 3.4.
Let $N$ be a hyperkähler manifold with complex structures $I,J,K$, and let
$\mathcal{J}_{I}=\left(\begin{array}[]{cc}I&0\\ 0&-I^{*}\end{array}\right),\quad\mathcal{J}_{J}=\left(\begin{array}[]{cc}J&0\\ 0&-J^{*}\end{array}\right),\quad\mathcal{J}_{K}=\left(\begin{array}[]{cc}K&0\\ 0&-K^{*}\end{array}\right)$
be the associated generalized complex structures on $\mathbb{T}N$. On $M=N\times N$, define
$\mathcal{I}_{1}=\mathcal{J}_{I}\oplus\mathcal{J}_{I},\quad\mathcal{I}_{2}=\mathcal{J}_{J}\oplus\mathcal{J}_{J},\quad\mathcal{I}_{3}=\mathcal{J}_{K}\oplus\left(-\mathcal{J}_{K}\right).$
Then each $\mathcal{I}_{a}$ is an integrable generalized complex structure, and
$\mathcal{I}_{a}\mathcal{I}_{b}+\mathcal{I}_{b}\mathcal{I}_{a}=-2\delta_{ab}{\rm Id}.$
However, $\mathcal{I}_{3}\neq\mathcal{I}_{1}\mathcal{I}_{2}.$ For any 2-form $B\in\Omega^{2}(M)$, set $H=dB$ and define
$\tilde{\mathcal{I}}_{a}=e^{B}\mathcal{I}_{a}e^{-B},\quad a=1,2,3.$
Then each $\tilde{\mathcal{I}}_{a}$ is integrable with respect to the $H$-twisted Courant bracket, the Clifford relations remain valid,
$\tilde{\mathcal{I}}_{a}\tilde{\mathcal{I}}_{b}+\tilde{\mathcal{I}}_{b}\tilde{\mathcal{I}}_{a}=-2\delta_{ab}\mathrm{Id},$
and
$\tilde{\mathcal{I}}_{3}\neq\tilde{\mathcal{I}}_{1}\tilde{\mathcal{I}}_{2}.$
4. The twistor space of a generalized Clifford manifold
Take
(4.1) $\mathcal{I}^{\pm}_{i}:=\frac{1}{2}\left(\mathcal{J}_{i}\pm\mathcal{I}_{i}\right)\ (i=1,2,3)\ \text{and}\ \mathcal{G}^{\pm}:=\frac{1}{2}({\rm Id}\pm\mathcal{G}).$
It is direct to check that
Lemma 4.1.
The endomorphisms $\mathcal{G}^{\pm}$ and $\mathcal{I}_{i}^{\pm}$ satisfy the following projection and quaternionic-type relations:
(4.2) $\begin{split}&(1)\ \left(\mathcal{G}^{\pm}\right)^{2}=\mathcal{G}^{\pm};\\ &(2)\ \mathcal{I}^{\pm}_{i}\mathcal{I}^{\pm}_{j}=-\delta_{ij}\mathcal{G}^{\pm}+\varepsilon_{ijk}\mathcal{I}^{\pm}_{k};\\ &(3)\ \mathcal{G}^{\pm}\mathcal{G}^{\mp}=\mathcal{G}^{\pm}\mathcal{I}^{\mp}_{i}=\mathcal{I}^{\pm}_{i}\mathcal{G}^{\mp}=\mathcal{I}^{\pm}_{i}\mathcal{I}^{\mp}_{j}=0.\end{split}$
Proof.
(1) Using $\mathcal{G}^{2}={\rm Id}$ we compute
$(\mathcal{G}^{\pm})^{2}=\frac{1}{4}({\rm Id}\pm\mathcal{G})^{2}=\frac{1}{4}({\rm Id}\pm 2\mathcal{G}+\mathcal{G}^{2})=\frac{1}{4}(2{\rm Id}\pm 2\mathcal{G})=\frac{1}{2}({\rm Id}\pm\mathcal{G})=\mathcal{G}^{\pm}.$
(2) By (1.2) and (4.1),
(4.3) $\begin{split}\mathcal{I}^{\pm}_{i}\mathcal{I}^{\pm}_{j}=&\frac{1}{4}\left(\mathcal{J}_{i}\pm\mathcal{I}_{i}\right)\left(\mathcal{J}_{j}\pm\mathcal{I}_{j}\right)\\ =&\frac{1}{4}\left(\mathcal{J}_{i}\mathcal{J}_{j}\pm\mathcal{J}_{i}\mathcal{I}_{j}\pm\mathcal{I}_{i}\mathcal{J}_{j}+\mathcal{I}_{i}\mathcal{I}_{j}\right)\\ =&\frac{1}{4}\left[-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{J}_{k}\pm(-\delta_{ij}\mathcal{G}+\epsilon_{ijk}\mathcal{I}_{k})\right.\\ &\left.\pm(-\delta_{ij}\mathcal{G}+\epsilon_{ijk}\mathcal{I}_{k})+(-\delta_{ij}{\rm Id}+\epsilon_{ijk}\mathcal{J}_{k})\right]\\ =&\frac{1}{2}\left[-\delta_{ij}\left({\rm Id\pm\mathcal{G}}\right)+\epsilon_{ijk}\left(\mathcal{J}_{k}\pm\mathcal{I}_{k}\right)\right]\\ =&-\delta_{ij}\mathcal{G}^{\pm}+\varepsilon_{ijk}\mathcal{I}^{\pm}_{k}.\end{split}$
(3) Similarly, (3) holds by (1.2) and (4.1). ∎
Corollary 4.1.
Let M be a Generalized Clifford manifold. If $\mathcal{G}$ is integrable, $\mathbb{T}M$ can split as
(4.4) $\mathbb{T}M=\mathbb{T}_{-1}\oplus\mathbb{T}_{1},$
where $\mathbb{T}_{-1}=\left\{A\in\mathbb{T}M|\mathcal{G}A=-A\right\}$ and $\mathbb{T}_{1}=\left\{A\in\mathbb{T}M|\mathcal{G}A=A\right\}$. Moreover, $\{\mathcal{I}^{-}_{1},\mathcal{I}^{-}_{2},\mathcal{I}^{-}_{3}\}$ is a hypercomplex structure over $\mathbb{T}_{-1}$ and $\{\mathcal{I}^{+}_{1},\mathcal{I}^{+}_{2},\mathcal{I}^{+}_{3}\}$ is a hypercomplex structure over $\mathbb{T}_{1}$.
We will use the decomposition to investigate the twistor space of Generalized Clifford manifold. Consider the stereographic projection,
$\zeta_{1}\mapsto\frac{1}{1+|\zeta_{1}|^{2}}\left(1-|\zeta_{1}|^{2},-\mathbf{i}(\zeta_{1}-\bar{\zeta_{1}}),-(\zeta_{1}+\bar{\zeta_{1}})\right),$
where $\zeta_{1}$ is a complex coordinate of a Riemann sphere. Take
(4.5) $T(\zeta_{1})=\left(\begin{array}[]{ccc}t_{11}&t_{12}&t_{13}\\ t_{21}&t_{22}&t_{23}\\ t_{31}&t_{32}&t_{33}\end{array}\right):=\frac{1}{1+|\zeta_{1}|^{2}}\left(\begin{array}[]{ccc}1-|\zeta_{1}|^{2}&-\mathbf{i}(\zeta_{1}-\bar{\zeta_{1}})&-(\zeta_{1}+\bar{\zeta_{1}})\\ \mathbf{i}(\zeta_{1}-\bar{\zeta_{1}})&1+\frac{1}{2}\left(\zeta_{1}^{2}+\bar{\zeta_{1}}^{2}\right)&-\frac{\mathbf{i}}{2}\left(\zeta_{1}^{2}-\bar{\zeta_{1}}^{2}\right)\\ \zeta_{1}+\bar{\zeta_{1}}&-\frac{\mathbf{i}}{2}\left(\zeta_{1}^{2}-\bar{\zeta_{1}}^{2}\right)&1-\frac{1}{2}\left(\zeta_{1}^{2}+\bar{\zeta_{1}}^{2}\right)\end{array}\right),$
which is a special orthogonal matrix. Consider the stereographic projection,
$\zeta_{2}\mapsto\frac{1}{1+|\zeta_{2}|^{2}}\left(1-|\zeta_{2}|^{2},-\mathbf{i}(\zeta_{2}-\bar{\zeta_{2}}),-(\zeta_{2}+\bar{\zeta_{2}})\right),$
where $\zeta_{2}$ is a complex coordinate of another Riemann sphere. Take
(4.6) $S(\zeta_{2})=\left(\begin{array}[]{ccc}s_{11}&s_{12}&s_{13}\\ s_{21}&s_{22}&s_{23}\\ s_{31}&s_{32}&s_{33}\end{array}\right):=\frac{1}{1+|\zeta_{2}|^{2}}\left(\begin{array}[]{ccc}1-|\zeta_{2}|^{2}&-\mathbf{i}(\zeta_{2}-\bar{\zeta_{2}})&-(\zeta_{2}+\bar{\zeta_{2}})\\ \mathbf{i}(\zeta_{2}-\bar{\zeta_{2}})&1+\frac{1}{2}\left(\zeta_{2}^{2}+\bar{\zeta_{2}}^{2}\right)&-\frac{\mathbf{i}}{2}\left(\zeta_{2}^{2}-\bar{\zeta_{2}}^{2}\right)\\ \zeta_{2}+\bar{\zeta_{2}}&-\frac{\mathbf{i}}{2}\left(\zeta_{2}^{2}-\bar{\zeta_{2}}^{2}\right)&1-\frac{1}{2}\left(\zeta_{2}^{2}+\bar{\zeta_{2}}^{2}\right)\end{array}\right),$
which is a special orthogonal matrix. Take
(4.7) $\begin{split}&\hat{\mathcal{I}}^{+}_{i}=t_{i1}\mathcal{I}^{+}_{1}+t_{i2}\mathcal{I}^{+}_{2}+t_{i3}\mathcal{I}^{+}_{3},\ (i=1,2,3),\\ &\hat{\mathcal{I}}^{-}_{i}=s_{i1}\mathcal{I}^{-}_{1}+s_{i2}\mathcal{I}^{-}_{2}+s_{i3}\mathcal{I}^{-}_{3},\ (i=1,2,3).\end{split}$
It is direct to check that
Lemma 4.2.
The quadruple $\left(\hat{\mathcal{I}}^{\pm}_{1},\hat{\mathcal{I}}^{\pm}_{2},\hat{\mathcal{I}}^{\pm}_{3},\mathcal{G}^{\pm}\right)$ satisfies the algebra of quaternion respectively, i.e.
  • (1)
[…]
  • (2)
[…]
Proof.
(1) Denote the $i$-th row vector of matrix $T$ as
$\tau_{i}=\left(t_{i1},t_{i2},t_{i3}\right),\quad i=1,2,3.$
It is direct to check that their vector product satisfy
(4.8) $\tau_{1}=\tau_{2}\times\tau_{3},\ \tau_{2}=\tau_{3}\times\tau_{1},\ \tau_{3}=\tau_{1}\times\tau_{2},$
i.e.
(4.9) $t_{1j}=t_{2r}t_{3s}\varepsilon_{rsj},\ t_{2j}=t_{3r}t_{1s}\varepsilon_{rsj},\ t_{3j}=t_{1r}t_{2s}\varepsilon_{rsj},\ j=1,2,3,$
where $\varepsilon_{ijk}$ is the Levi-Civita symbol
[…]
By the orthogonal property of $T$ and (4.9), we have
[…]
(2) Obviously, conclusion (2) holds by the third equation (4.2). ∎
Take
(4.11) $\mathcal{K}_{i}(\zeta_{1},\zeta_{2}):=\hat{\mathcal{I}}^{+}_{i}+\hat{\mathcal{I}}^{-}_{i}=\sum_{l=1}^{3}\left(t_{il}\mathcal{I}_{l}^{+}+s_{il}\mathcal{I}_{l}^{-}\right),\ i=1,2,3.$
For simplicity, we denote $\mathcal{K}_{i}(\zeta_{1},\zeta_{2})$ by $\mathcal{K}_{i}$. Then by (4.1),(4.5), (4.6) and (4.7), (4.11) can be rewritten as
$\left(\begin{array}[]{c}\mathcal{K}_{1}\\ \mathcal{K}_{2}\\ \mathcal{K}_{3}\end{array}\right)=\frac{1}{2}T\left(\begin{array}[]{c}\mathcal{I}_{1}+\mathcal{I}_{2}\mathcal{I}_{3}\\ \mathcal{I}_{2}+\mathcal{I}_{3}\mathcal{I}_{1}\\ \mathcal{I}_{3}+\mathcal{I}_{1}\mathcal{I}_{2}\end{array}\right)+\frac{1}{2}S\left(\begin{array}[]{c}-\mathcal{I}_{1}+\mathcal{I}_{2}\mathcal{I}_{3}\\ -\mathcal{I}_{2}+\mathcal{I}_{3}\mathcal{I}_{1}\\ -\mathcal{I}_{3}+\mathcal{I}_{1}\mathcal{I}_{2}\end{array}\right).$
Now we can prove Theorem 1.2.
Proof of Theorem 1.2 By the definition of $\mathcal{K}_{i}$ in (4.11) and Lemma 4.2, one checks directly that
(4.12) $\mathcal{K}_{i}^{2}=-\operatorname{Id}(i=1,2,3),\quad\mathcal{K}_{i}\mathcal{K}_{j}=-\mathcal{K}_{j}\mathcal{K}_{i}(i\neq j),$
and
(4.13) $\left\langle\mathcal{K}_{i}\cdot,\mathcal{K}_{i}\cdot\right\rangle=\langle\cdot,\cdot\rangle.$
Hence $\left(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}\right)$ forms an almost generalized Clifford structure.
We now prove that these three structures $\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}$ are simultaneously integrable. Define
$\alpha_{i}^{l}:=\frac{1}{2}\left(t_{il}+s_{il}\right),\quad\beta_{i}^{l}:=\frac{1}{2}\left(t_{il}-s_{il}\right).$
With this notation, $\mathcal{K}_{i}$ can be written as the linear combination
$\mathcal{K}_{i}=\sum_{l=1}^{3}\left(\alpha_{i}^{l}\mathcal{I}_{l}+\beta_{i}^{l}\mathcal{J}_{l}\right).$
By the bilinearity of Nijenhuis tensor, we have
(4.14) $\mathcal{N}\left(\mathcal{K}_{i},\mathcal{K}_{j}\right)=\sum_{l,m=1}^{3}\left(\alpha_{i}^{l}\alpha_{j}^{m}\mathcal{N}\left(\mathcal{I}_{l},\mathcal{I}_{m}\right)+\alpha_{i}^{l}\beta_{j}^{m}\mathcal{N}\left(\mathcal{I}_{l},\mathcal{J}_{m}\right)+\beta_{i}^{l}\alpha_{j}^{m}\mathcal{N}\left(\mathcal{J}_{l},\mathcal{I}_{m}\right)+\beta_{i}^{l}\beta_{j}^{m}\mathcal{N}\left(\mathcal{J}_{l},\mathcal{J}_{m}\right)\right).$
Then, by Theorem 1.1, we obtain
(4.15) $\mathcal{N}(\mathcal{K}_{i},\mathcal{K}_{j})=0,\ i,j=1,2,3,$
i.e., $\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}$ are simultaneously integrable. Then by (4.12)-(4.15), $\left(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}\right)$ forms a generalized Clifford structure. ∎
Before proving Theorem 1.3, we briefly explain the strategy. The generalized almost complex structure on the twistor space splits into the generalized complex structure $\hat{I}(\zeta_{1},\zeta_{2})$ along $M$ and the canonical complex structure $\mathcal{J}$ along $S^{2}\times S^{2}$. Accordingly, the integrability reduces to checking the closure of the eigenbundle under the Dorfman bracket in four cases: the $M$-$M$ term, the $(S^{2}\times S^{2})$-$(S^{2}\times S^{2})$ term, and the two mixed terms. The first two follow from the integrability of $\hat{I}(\zeta_{1},\zeta_{2})$ and $\mathcal{J}$, while the mixed terms are handled by Lemmas 4.3–4.6.
Proof of Theorem 1.3. Let $Z:=M\times S$ with $S:=S^{2}_{\zeta_{1}}\times S^{2}_{\zeta_{2}}\simeq\mathbb{CP}^{1}\times\mathbb{CP}^{1}$. Write
$\mathbb{T}Z:=TZ\oplus T^{*}Z,\qquad\mathbb{T}M:=TM\oplus T^{*}M,\qquad\mathbb{T}S:=TS\oplus T^{*}S,$
so that $\mathbb{T}Z\simeq\mathbb{T}M\oplus\mathbb{T}S$. We use the Dorfman bracket on $\mathbb{T}Z$:
(4.16) $X+\alpha,\;Y+\beta]_{D}=[X,Y]+\mathcal{L}_{X}\beta-\iota_{Y}d\alpha,$
where $X,Y\in\Gamma(TZ)$, $\alpha,\beta\in\Omega^{1}(Z)$, and the anchor is $\rho(X+\alpha)=X$.
Let $J_{S}$ denote the standard complex structure on $S\simeq\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, and let
(4.17) $\mathcal{J}:=\begin{pmatrix}-J_{S}&0\\ 0&J_{S}^{T}\end{pmatrix}\in\Gamma(\mathrm{End}(\mathbb{T}S))$
be the induced generalized complex structure on $\mathbb{T}S$. By (4.11) and Theorem 1.2, we have a family of generalized complex structures over $\mathbb{T}M$ by
(4.18) $\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})=\vec{c}(\zeta_{1})\cdot\vec{\mathcal{J}}^{(+)}+\vec{d}(\zeta_{2})\cdot\vec{\mathcal{J}}^{(-)},$
where $\vec{\mathcal{J}}^{(\pm)}:=(\mathcal{J}_{1}^{(\pm)},\mathcal{J}_{2}^{(\pm)},\mathcal{J}_{3}^{(\pm)})$ and
(4.19) $\vec{c}(\zeta_{1})=\left(\frac{1-|\zeta_{1}|^{2}}{1+|\zeta_{1}|^{2}},\ \frac{\mathbf{i}(\bar{\zeta}_{1}-\zeta_{1})}{1+|\zeta_{1}|^{2}},\ -\frac{\zeta_{1}+\bar{\zeta}_{1}}{1+|\zeta_{1}|^{2}}\right),\quad\vec{d}(\zeta_{2})=\left(\frac{1-|\zeta_{2}|^{2}}{1+|\zeta_{2}|^{2}},\ \frac{\mathbf{i}(\bar{\zeta}_{2}-\zeta_{2})}{1+|\zeta_{2}|^{2}},\ -\frac{\zeta_{2}+\bar{\zeta}_{2}}{1+|\zeta_{2}|^{2}}\right).$
Finally, define
(4.20) $\hat{\mathbb{I}}:=\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})\ \oplus\ \mathcal{J}\ \in\ \Gamma(\mathrm{End}(\mathbb{T}Z)).$
For an almost generalized complex structure $\mathbb{I}$, let $L_{\mathbb{I}}\subset(\mathbb{T}Z)\otimes\mathbb{C}$ be the $\mathbf{i}$–eigenbundle. Since $\hat{\mathbb{I}}$ is block diagonal with respect to $\mathbb{T}Z\simeq\mathbb{T}M\oplus\mathbb{T}S$, we have
(4.21) $L_{\hat{\mathbb{I}}}=L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})}\ \oplus\ L_{\mathcal{J}}.$
Moreover, because $\mathcal{J}$ is induced by the complex structure $J_{S}$,
(4.22) $L_{\mathcal{J}}=T^{0,1}S\ \oplus\ T^{*1,0}S.$
In local holomorphic coordinates $(\zeta_{1},\zeta_{2})$ on $S$, this means that $L_{\mathcal{J}}$ is locally generated over $\mathbb{C}$ by
(4.23) $\partial_{\bar{\zeta}_{1}},\quad\partial_{\bar{\zeta}_{2}},\quad d\zeta_{1},\quad d\zeta_{2}.$
Integrability of $\hat{\mathbb{I}}$ is equivalent to the involutivity of its +i+i–eigenbundle:
[…]
(4.25) $u+\alpha,\ v+\beta]_{D}=[u,v]_{D}+[\alpha,\beta]_{D}+[\alpha,v]_{D}+[u,\beta]_{D}.$
We will show that each term in (4.25) lands in $\Gamma(L_{\hat{\mathcal{I}}})\oplus\Gamma(L_{\mathcal{J}})$.
(1) The SSSS term: $\Gamma(L_{\mathcal{J}}),\Gamma(L_{\mathcal{J}})]_{D}$. Since $J_{S}$ is integrable, the induced generalized complex structure $\mathcal{J}$ on $\mathbb{T}S$ is integrable, hence
(4.26) $\Gamma(L_{\mathcal{J}}),\Gamma(L_{\mathcal{J}})]_{D}\subset\Gamma(L_{\mathcal{J}}).$
(2) The MMMM term: $\Gamma(L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})}),\Gamma(L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})})]_{D}$. By Theorem 1.2, for each fixed $(\zeta_{1},\zeta_{2})\in S$, the generalized almost complex structure $\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})$ on $M$ is integrable with respect to the Dorfman bracket on $\mathbb{T}M$, i.e.,
(4.27) $\Gamma(L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})}),\Gamma(L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})})]_{D}\subset\Gamma(L_{\hat{\mathcal{I}}(\zeta_{1},\zeta_{2})}).$
(3) The SSMM term: $\Gamma(L_{\mathcal{J}}),\Gamma(L_{\hat{\mathcal{I}}})]_{D}$. We first record a precise mixed bracket identity.
Lemma 4.3.
Let $\alpha=Y+\eta\in\Gamma(\mathbb{T}S)$ and $v=X+\xi\in\Gamma(\mathbb{T}M)$, viewed as sections of $\mathbb{T}Z$. Then
(4.28) $\alpha,v]_{D}=\mathcal{L}_{Y}v=\mathcal{L}_{\rho(\alpha)}v.$
Proof.
By (4.16),
$Y+\eta,\ X+\xi]_{D}=[Y,X]+\mathcal{L}_{Y}\xi-\iota_{X}d\eta.$
Here $Y,X]=\mathcal{L}_{Y}X$ because $Y$ differentiates only the coefficients of $X$ (which may depend on the SS–coordinates). Moreover, $d\eta$ has only SS–components while $X$ is tangent to $M$, hence $\iota_{X}d\eta=0$. Therefore
$Y+\eta,\ X+\xi]_{D}=\mathcal{L}_{Y}X+\mathcal{L}_{Y}\xi=\mathcal{L}_{Y}(X+\xi),$
which is (4.28). ∎
[…]
Since $v$ may depend on $(\zeta_{1},\zeta_{2})$, (4.29) is not automatic and is the genuinely nontrivial part of the argument.
Define $\mathbb{R}^{3}$–valued 11–forms on $S$ by
(4.30) $\boldsymbol{\omega}_{\zeta_{1}}:=\vec{c}\times{\rm d}\vec{c},\qquad\boldsymbol{\omega}_{\zeta_{2}}:=\vec{d}\times{\rm d}\vec{d},$
and an $\mathrm{End}(\mathbb{T}M)$–valued 11–form by
(4.31) $\Omega:=\boldsymbol{\omega}_{\zeta_{1}}\cdot\vec{\mathcal{J}}^{(+)}+\boldsymbol{\omega}_{\zeta_{2}}\cdot\vec{\mathcal{J}}^{(-)}.$
Lemma 4.4.
The family $\hat{\mathcal{I}}$ satisfies
(4.32) ${\rm d}\hat{\mathcal{I}}=\frac{1}{2}[\Omega,\hat{\mathcal{I}}$],
where $d$ denotes the exterior derivative along the parameter space $S$. Hence, taking (0,1)(0,1)–parts,
(4.33) $\bar{\partial}_{S}\hat{\mathcal{I}}=[V^{0,1},\hat{\mathcal{I}}],\qquad V^{0,1}:=\frac{1}{2}\,\Omega^{0,1}.$
Proof.
We use two elementary identities.
(i) Cross-product commutator rule. For any $\vec{a},\vec{b}\in\mathbb{R}^{3}$, using Lemma 4.2 one computes
(4.34) $\vec{a}\cdot\vec{\mathcal{J}}^{(\pm)},\ \vec{b}\cdot\vec{\mathcal{J}}^{(\pm)}]=2(\vec{a}\times\vec{b})\cdot\vec{\mathcal{J}}^{(\pm)},$
and
$\vec{a}\cdot\vec{\mathcal{J}}^{(+)},\ \vec{b}\cdot\vec{\mathcal{J}}^{(-)}]=0.$
(ii) Unit vector identity. Since $|\vec{c}|^{2}=1$, we have $\vec{c}\cdot d\vec{c}=0$. With $\boldsymbol{\omega}_{\zeta_{1}}=\vec{c}\times d\vec{c}$, the triple product identity $(a\times b)\times a=(a\cdot a)b-(a\cdot b)a$ yields
(4.35) $\boldsymbol{\omega}_{\zeta_{1}}\times\vec{c}=(\vec{c}\times{\rm d}\vec{c})\times\vec{c}={\rm d}\vec{c}.$
Similarly, $\boldsymbol{\omega}_{\zeta_{2}}\times\vec{d}={\rm d}\vec{d}$.
Now differentiate (4.18):
${\rm d}\hat{\mathcal{I}}=({\rm d}\vec{c})\cdot\vec{\mathcal{J}}^{(+)}+({\rm d}\vec{d})\cdot\vec{\mathcal{J}}^{(-)}.$
Using (4.35) and (4.34) we obtain
$({\rm d}\vec{c})\cdot\vec{\mathcal{J}}^{(+)}=(\boldsymbol{\omega}_{\zeta_{1}}\times\vec{c})\cdot\vec{\mathcal{J}}^{(+)}=\frac{1}{2}[\boldsymbol{\omega}_{\zeta_{1}}\cdot\vec{\mathcal{J}}^{(+)},\ \vec{c}\cdot\vec{\mathcal{J}}^{(+)}].$
An analogous formula holds for the ()(-)–part. Since the $(+)$ and ()(-) sectors commute, combining both parts gives
${\rm d}\hat{\mathcal{I}}=\frac{1}{2}[\Omega,\hat{\mathcal{I}}$],
which is (4.32). Taking (0,1)(0,1)–parts yields (4.33). ∎
Define a (0,1)(0,1)–operator on sections of $(\mathbb{T}M)\otimes\mathbb{C}$ along $S$ by
(4.36) $\nabla^{0,1}_{Y}v:=\mathcal{L}_{Y}v-V^{0,1}(Y)\,v,\qquad Y\in T^{0,1}S.$
Lemma 4.5.
[…]
Equivalently, $L_{\hat{\mathcal{I}}}$ is preserved by $\nabla^{0,1}$.
Proof.
From $\hat{\mathcal{I}}v=\mathbf{i}v$ we take Lie derivative along $Y$ and use Leibniz:
(4.38) $(\mathcal{L}_{Y}\hat{\mathcal{I}})v+\hat{\mathcal{I}}(\mathcal{L}_{Y}v)=\mathbf{i}\,\mathcal{L}_{Y}v.$
By Lemma 4.4, $\mathcal{L}_{Y}\hat{\mathcal{I}}=[V^{0,1}(Y),\hat{\mathcal{I}}$]. Substituting into (4.38) gives
$V^{0,1}(Y),\hat{\mathcal{I}}]v+(\hat{\mathcal{I}}-\mathbf{i})\mathcal{L}_{Y}v=0.$
Since $\hat{\mathcal{I}}v=\mathbf{i}v$, we have
$V^{0,1}(Y),\hat{\mathcal{I}}]v=V^{0,1}(Y)(\hat{\mathcal{I}}v)-\hat{\mathcal{I}}(V^{0,1}(Y)v)=\mathbf{i}\,V^{0,1}(Y)v-\hat{\mathcal{I}}(V^{0,1}(Y)v)=-(\hat{\mathcal{I}}-\mathbf{i})\bigl(V^{0,1}(Y)v\bigr).$
Hence
$(\hat{\mathcal{I}}-\mathbf{i})\bigl(\mathcal{L}_{Y}v-V^{0,1}(Y)v\bigr)=0,$
which is exactly (4.37) using (4.36). ∎
We now show that $\nabla^{0,1}$ is flat on the parameter space $S$, so that we can choose local frames $v_{a}$ with $\nabla^{0,1}v_{a}=0$.
Lemma 4.6.
The (0,1)(0,1)–connection $\nabla^{0,1}=\bar{\partial}_{S}-V^{0,1}$ is flat:
(4.39) $(\nabla^{0,1})^{2}=0.$
Proof.
The curvature of $\nabla^{0,1}$ is the (0,2)(0,2)–form with values in $\mathrm{End}(\mathbb{T}M\otimes\mathbb{C})$ given by the Maurer–Cartan expression
$(\nabla^{0,1})^{2}=-\bigl(\bar{\partial}_{S}V^{0,1}-V^{0,1}\wedge V^{0,1}\bigr).$
We show $\bar{\partial}_{S}V^{0,1}=0$ and $V^{0,1}\wedge V^{0,1}=0$.
Decompose
$V^{0,1}=V_{\zeta_{1}}^{0,1}+V_{\zeta_{2}}^{0,1},$
where $V_{\zeta_{1}}^{0,1}$ depends solely on $\left(\zeta_{1},\bar{\zeta}_{1}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{1}$, whereas $V_{\zeta_{2}}^{0,1}$ depends solely on $\left(\zeta_{2},\bar{\zeta}_{2}\right)$ and is proportional to $\mathrm{d}\bar{\zeta}_{2}$. This follows from the definition $V^{0,1}=\tfrac{1}{2}\Omega^{0,1}$ and from (4.30) because $\vec{c}$ depends only on $\zeta_{1}$ and $\vec{d}$ only on $\zeta_{2}$. Hence $V^{0,1}_{\zeta_{1}}=A(\zeta_{1},\bar{\zeta}_{1})\,{\rm d}\bar{\zeta}_{1}$ and $V^{0,1}_{\zeta_{2}}=B(\zeta_{2},\bar{\zeta}_{2})\,{\rm d}\bar{\zeta}_{2}$, where
$A(\zeta_{1},\bar{\zeta}_{1})\in\mathrm{span}\{\mathcal{J}_{i}^{(+)}\},\qquad B(\zeta_{2},\bar{\zeta}_{2})\in\mathrm{span}\{\mathcal{J}_{i}^{(-)}\}.$
Since $d\bar{\zeta}_{1}\wedge d\bar{\zeta}_{1}=0$ and $d\bar{\zeta}_{2}\wedge d\bar{\zeta}_{2}=0$, we have $\bar{\partial}_{S}V^{0,1}_{\zeta_{1}}=0$ and $\bar{\partial}_{S}V^{0,1}_{\zeta_{2}}=0$, thus $\bar{\partial}_{S}V^{0,1}=0$.
For the wedge term, note first that $V^{0,1}_{\zeta_{1}}\wedge V^{0,1}_{\zeta_{1}}=0$ and $V^{0,1}_{\zeta_{2}}\wedge V^{0,1}_{\zeta_{2}}=0$ because each is a (0,1)(0,1)–form proportional to a single generator. For the mixed wedge, using graded antisymmetry,
$V^{0,1}_{\zeta_{1}}\wedge V^{0,1}_{\zeta_{2}}+V^{0,1}_{\zeta_{2}}\wedge V^{0,1}_{\zeta_{1}}=[A,B]\ d\bar{\zeta_{1}}\wedge d\bar{\zeta_{2}}.$
But $[A,B]=0$ by (4.2) since $A$ is valued in the $(+)$ sector and $B$ in the ()(-) sector. Therefore $V^{0,1}\wedge V^{0,1}=0$. This proves $(\nabla^{0,1})^{2}=0$. ∎
Corollary 4.2.
Locally on $S$, there exist sections $\{v_{a}\}$ spanning $L_{\hat{\mathcal{I}}}$ such that
(4.40) $\nabla^{0,1}v_{a}=0\quad\Longleftrightarrow\quad\mathcal{L}_{Y}v_{a}=V^{0,1}(Y)\,v_{a}\quad\forall\,Y\in T^{0,1}S.$
Proof.
By Lemma 4.6, $\nabla^{0,1}$ defines a flat $\bar{\partial}$–operator along $S$. A standard local trivialization argument for flat (0,1)(0,1)–connections yields local frames with $\nabla^{0,1}v_{a}=0$. Lemma 4.5 ensures these frames may be chosen inside $L_{\hat{\mathcal{I}}}$ and hence span it. ∎
Let $\alpha\in\Gamma(L_{\mathcal{J}})$ and let $v_{a}$ be as in Corollary 4.2. There are two cases:
  • (1)
[…]
  • (2)
    If $\alpha\in T^{0,1}S$ (e.g. $\alpha=\partial_{\bar{\zeta}_{1}}$ or $\partial_{\bar{\zeta}_{2}}$), then $\rho(\alpha)=\alpha=:Y$ and Lemma 4.3 together with (4.40) gives
[…]
Thus
[…]
$u,\beta]_{D}+[\beta,u]_{D}=\mathcal{D}\langle u,\beta\rangle,$
where $\mathcal{D}$ is the Courant algebroid differential. Because $\mathbb{T}Z=\mathbb{T}M\oplus\mathbb{T}S$ is an orthogonal direct sum with respect to the Courant pairing, any $u\in\Gamma(\mathbb{T}M)$ and $\beta\in\Gamma(\mathbb{T}S)$ satisfy $\langle u,\beta\rangle=0$, and hence $\mathcal{D}\langle u,\beta\rangle=0$. So we have
[…]
by(4.41). A symmetric argument (or a direct inspection from (4.16) using that $\rho(u)$ is tangent to $M$ while $\beta$ is tangent to $S$) shows that
[…]
5. T-duality and rank-3 generalized Clifford structures
The formulation of T-duality used here follows the generalized-geometric approach of Cavalcanti and Gualtieri [5], building on the topological T-duality framework for circle and torus bundles with $H$-flux developed in [2, 3]. In this section, we show that rank-3 twisted generalized Clifford structures are stable under T-duality. More precisely, we prove that T-duality preserves not only the original Clifford triple $(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})$, but also the induced structures $(\mathcal{J}_{1},\mathcal{J}_{2},\mathcal{J}_{3},\mathcal{G})$ and the associated Spin(3)-rotated family $(\mathcal{K}_{1}(\zeta_{1},\zeta_{2}),\mathcal{K}_{2}(\zeta_{1},\zeta_{2}),\mathcal{K}_{3}(\zeta_{1},\zeta_{2}))$. This shows that the Clifford-to-twistor construction developed in the previous sections is compatible with the natural symmetry of generalized geometry given by T-duality.
[…]
$\Phi:\mathbb{T}M=TM\oplus T^{*}M\longrightarrow\mathbb{T}\widetilde{M}=T\widetilde{M}\oplus T^{*}\widetilde{M}$
such that
(5.1) $\langle\Phi(A),\Phi(B)\rangle=\langle A,B\rangle,$
and
(5.2) $\Phi\bigl([A,B]_{H}\bigr)=[\Phi(A),\Phi(B)]_{\widetilde{H}},\qquad A,B\in\Gamma(\mathbb{T}M).$
Here $\cdot,\cdot]_{H}$ and $\cdot,\cdot]_{\widetilde{H}}$ denote the $H$-twisted and $\widetilde{H}-$twisted Dorfman brackets, respectively.
Condition (5.2) implies that any generalized-geometric structure defined in terms of the natural pairing and the twisted Dorfman bracket can be transported through $\Phi$. In particular, for any bundle endomorphism $\mathcal{I}\in\Gamma(End(\mathbb{T}M))$, we define its T-dual transform by
$).\widetilde{\mathcal{I}}=\Phi\mathcal{I}\Phi^{-1}\in\Gamma(End(\mathbb{T}\widetilde{M}$)).
Lemma 5.1.
Let $\mathcal{I},\mathcal{J}\in\Gamma(End(\mathbb{T}M))$. Then the mixed generalized Nijenhuis tensor is preserved by T-duality in the sense that
$\mathcal{N}_{\widetilde{H}}(\widetilde{\mathcal{I}},\widetilde{\mathcal{J}})(\Phi A,\Phi B)=\Phi\Bigl(\mathcal{N}_{H}(\mathcal{I},\mathcal{J})(A,B)\Bigr),\qquad A,B\in\Gamma(\mathbb{T}M).$
In particular,
$\mathcal{N}_{H}(\mathcal{I},\mathcal{J})=0\quad\Longrightarrow\quad\mathcal{N}_{\widetilde{H}}(\widetilde{\mathcal{I}},\widetilde{\mathcal{J}})=0.$
Proof.
This follows directly from the definition of the generalized Nijenhuis concomitant and the fact that $\Phi$ intertwines the twisted Dorfman brackets:
$\Phi([A,B]_{H})=[\Phi(A),\Phi(B)]_{\widetilde{H}}$.
Indeed, every term in the definition of $\mathcal{N}(\mathcal{I},\mathcal{J})$ is functorial under conjugation by $\Phi$, and therefore the whole expression is transported to the corresponding one on $\mathbb{T}\widetilde{M}$. ∎
We now show that T-duality preserves the class of rank-3 twisted generalized Clifford manifolds.
Proposition 5.1.
[…]
be an orthogonal Courant algebroid isomorphism. Assume that $(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})$ is an $H$-twisted rank-3 generalized Clifford structure on $M$. Define
$\widetilde{\mathcal{I}}_{a}:=\Phi\mathcal{I}_{a}\Phi^{-1},\qquad a=1,2,3.$
Then $(\widetilde{I}_{1},\widetilde{I}_{2},\widetilde{I}_{3})$ is a $\widetilde{H}-$twisted rank-3 generalized Clifford structure on $\widetilde{M}$.
Proof.
Since each $\mathcal{I}_{a}$ is orthogonal and satisfies $\mathcal{I}_{a}^{2}=-{\rm Id}$, and since $\Phi$ is orthogonal, each $\widetilde{I}_{a}$ is again an orthogonal almost generalized complex structure on $\mathbb{T}\widetilde{M}$:
$\widetilde{\mathcal{I}}_{a}^{2}=\Phi\mathcal{I}_{a}^{2}\Phi^{-1}=-{\rm Id}.$
Moreover, the Clifford relations are preserved:
$\widetilde{\mathcal{I}}_{a}\widetilde{\mathcal{I}}_{b}+\widetilde{\mathcal{I}}_{b}\widetilde{\mathcal{I}}_{a}=\Phi(\mathcal{I}_{a}\mathcal{I}_{b}+\mathcal{I}_{b}\mathcal{I}_{a})\Phi^{-1}=-2\delta_{ab}{\rm Id}.$
Finally, since each $I_{a}$ is integrable with respect to the $H$-twisted Dorfman bracket, Lemma 5.1 gives
$\mathcal{N}_{\widetilde{H}}(\widetilde{I}_{a},\widetilde{I}_{a})=0,\qquad a=1,2,3.$
Hence $(\widetilde{I}_{1},\widetilde{I}_{2},\widetilde{I}_{3})$ defines a $\widetilde{H}-$twisted rank-3 generalized Clifford structure on $\widetilde{M}$. ∎
We now show that these induced structures (1.1) are also compatible with T-duality.
Proposition 5.2.
With the notation (1.1), define
(5.3) $\widetilde{\mathcal{J}}_{i}:=\frac{1}{2}\varepsilon_{ijk}\widetilde{\mathcal{I}}_{j}\widetilde{\mathcal{I}}_{k},\qquad\widetilde{\mathcal{G}}:=-\widetilde{\mathcal{I}}_{1}\widetilde{\mathcal{I}}_{2}\widetilde{\mathcal{I}}_{3}.$
Then
$\widetilde{\mathcal{J}}_{i}=\Phi\mathcal{J}_{i}\Phi^{-1},\qquad\widetilde{\mathcal{G}}=\Phi\mathcal{G}\Phi^{-1}.$
Proof.
Using the definition of $\mathcal{J}_{i}$ in (1.1) and the multiplicativity of conjugation, we obtain
$\widetilde{\mathcal{J}}_{i}=\frac{1}{2}\varepsilon_{ijk}\widetilde{\mathcal{I}}_{j}\widetilde{\mathcal{I}}_{k}=\frac{1}{2}\varepsilon_{ijk}(\Phi\mathcal{I}_{j}\Phi^{-1})(\Phi\mathcal{I}_{k}\Phi^{-1})=\Phi\Bigl(\frac{1}{2}\varepsilon_{ijk}\mathcal{I}_{j}\mathcal{I}_{k}\Bigr)\Phi^{-1}=\Phi\mathcal{J}_{i}\Phi^{-1}.$
Similarly,
$\widetilde{\mathcal{G}}=-\widetilde{\mathcal{I}}_{1}\widetilde{\mathcal{I}}_{2}\widetilde{\mathcal{I}}_{3}=-\Phi\mathcal{I}_{1}\mathcal{I}_{2}\mathcal{I}_{3}\Phi^{-1}=\Phi\mathcal{G}\Phi^{-1}.$
As an immediate consequence, all algebraic identities in (1.2) are preserved by T-duality. Define similarly
[…]
and let
$\bigl(\widetilde{\mathcal{K}}_{1}(\zeta_{1},\zeta_{2}),\widetilde{\mathcal{K}}_{2}(\zeta_{1},\zeta_{2}),\widetilde{\mathcal{K}}_{3}(\zeta_{1},\zeta_{2})\bigr)$
be the corresponding Clifford-rotated family on $\widetilde{M}$.
Proposition 5.3.
For every $(\zeta_{1},\zeta_{2})\in\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, one has
$\widetilde{\mathcal{K}}_{i}(\zeta_{1},\zeta_{2})=\Phi{\mathcal{K}}_{i}(\zeta_{1},\zeta_{2})\Phi^{-1},\qquad i=1,2,3.$
Equivalently, T-duality commutes with the rank-3 Clifford rotations.
Proof.
By Proposition 5.2, we have
$\widetilde{\mathcal{I}}_{i}^{\pm}=\frac{1}{2}(\widetilde{\mathcal{J}}_{i}\pm\widetilde{\mathcal{I}}_{i})=\frac{1}{2}\Phi({\mathcal{J}}_{i}\pm{\mathcal{I}}_{i})\Phi^{-1}=\Phi{\mathcal{I}}_{i}^{\pm}\Phi^{-1}.$
Hence, using the definition of the rotated structures in Section 4,
$\widehat{\widetilde{\mathcal{I}}}_{i}^{+}=\sum_{l=1}^{3}t_{il}\widetilde{\mathcal{I}}_{l}^{+}=\sum_{l=1}^{3}t_{il}\Phi I_{l}^{+}\Phi^{-1}=\Phi\Bigl(\sum_{l=1}^{3}t_{il}{\mathcal{I}}_{l}^{+}\Bigr)\Phi^{-1}=\Phi\widehat{\mathcal{I}}_{i}^{\,+}\Phi^{-1},$
and similarly
$\widehat{\widetilde{\mathcal{I}}}_{i}^{-}=\Phi\widehat{\mathcal{I}}_{i}^{\,-}\Phi^{-1},$
Therefore,
$\widetilde{\mathcal{K}}_{i}=\frac{1}{2}\bigl(\widetilde{\widehat{I}}_{i}^{\,+}+\widetilde{\widehat{I}}_{i}^{\,-}\bigr)=\Phi\Bigl(\frac{1}{2}(\widehat{{\mathcal{I}}}_{i}^{+}+\widehat{{\mathcal{I}}}_{i}^{-})\Bigr)\Phi^{-1}=\Phi\mathcal{K}_{i}\Phi^{-1}.$
This proves the claim. ∎
Combining Theorem 5.1 with Proposition 5.3, we obtain:
Corollary 5.1.
The T-dual of the Spin(3)-rotated family associated with a rank-3 twisted generalized Clifford structure is exactly the Spin(3)-rotated family associated with the T-dual Clifford triple. In particular, the construction of the $S^{2}\times S^{2}$-family of generalized complex structures is natural with respect to T-duality.
Remark 5.1.
Proposition 5.1 and Proposition 5.3 show that the $M$-component of the twistor generalized complex structure constructed in Theorem 1.3 is compatible with T-duality. It is therefore natural to expect that, after pulling back the twist to $M\times S^{2}\times S^{2}$, the corresponding generalized complex structure on the twistor space admits a T-dual interpretation as well. We leave this total-space formulation to future work.